OpenStudy (anonymous):

If \[x ^{2}-xy+y ^{2} = a\] and \[x ^{3} + y ^{3} = b\] then xy = ?

OpenStudy (experimentx):

use a^3+b^3 formula to find out the value of x+y first.

OpenStudy (anonymous):

\[x ^{3} + y^{3} = (x + y)(x^{2}-xy+y^{2})\] b = (x +y) (a) b/a = x + y

OpenStudy (anonymous):

\[xy=\frac{b^2-a^4}{3a^2}\]

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

\[xy=\frac{b^2-a^3}{3a}\]

OpenStudy (experimentx):

square x+y put the values of x^2+y^2 from your first equation.

OpenStudy (anonymous):

sorry it is wrong again..

OpenStudy (anonymous):

\[xy=\frac{b^2-a^3}{3a^2}\]

OpenStudy (anonymous):

take the sq of last eq. you found..

OpenStudy (anonymous):

so its x^2 + y^2 = b^2/a^2 - 2x .. im trying lol

OpenStudy (anonymous):

\[\frac{b^2}{a^2}=x^2+2xy+y^2\]

OpenStudy (anonymous):

wait ill try to see my answer then ill compare it to cinar

OpenStudy (anonymous):

\[x^2+y^2=a+xy\]

OpenStudy (anonymous):

\[\frac{b^2}{a^2}=a+3xy\] \[\frac{b^2}{a^2}-a=3xy\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw